# Decentralized Aerial Manipulation of a Cable-Suspended Load using Multi-Agent Reinforcement Learning

### v2

@@ 1 Introduction
Autonomous Micro Aerial Vehicles (MAVs) offer great capability for transporting slung loads to dangerous and remote locations [1]. While a single low-cost MAV has limited payload capacity, collaborative teams of MAVs can transport significantly heavier loads. In addition, by connecting each MAV with the load at different points using tethers, the full pose of the load can be controlled by changing the position of the MAVs, yielding a cooperative cable-suspended manipulation solution, which shows great potential for aerial-based construction, inspection, and resecuring [2, 3, 4, 5, 6].
@@
To coordinate and control MAV fleets, the state-of-the-art method [6] employs a centralized framework that accurately captures the strong dynamical coupling between the MAVs and the suspended load. This ensures safety and stability while addressing the significant underactuation inherent to cable-suspended systems, preventing actuator saturations and reciprocal collisions. However, using centralized control strategies for such systems suffers from critical drawbacks: computational complexity tends to scale exponentially with the number of agents for many approaches, rendering real-time control infeasible for larger teams with a centralized scheme [6, 7]. In addition, dependence on global state information and centralized communication is often impractical due to limits on sensors and communication bandwidth. A plausible solution, decentralization, remains an open challenge to effectively coordinate MAV fleets due to partial observability, limited communication bandwidth, and decision-making under strong dynamical coupling between agents while co-manipulating an object.
@@
In this work, we present the first decentralized algorithm to achieve a real-world demonstrated full-pose manipulation of a cable-suspended payload using a team of MAVs. Our method leverages multi-agent reinforcement learning (MARL) and does not require any inter-agent communication. Instead, each agent only takes their own state and identity, the load pose, and the target load pose as observations. We train the policy through MARL in a centralized training with decentralized execution (CTDE) paradigm using multi-agent proximal policy optimization (MAPPO) [8]. Each MAV learns to communicate implicitly through the load pose information. To fill the sim-to-real gap in this highly dynamic cooperative task, we design the action space of the reinforcement learning (RL) policy as reference linear accelerations and body rates of the MAV and combine the RL policy with a low-level controller based on incremental nonlinear dynamic inversion (INDI) [9, 10, 11]. The low-level controller follows the linear acceleration command with the body rate reference as the feedforward commands, ensuring agile and smooth control maneuvers during the cooperative manipulation.
@@
Our method enables zero-shot transfer of the policy from simulation to real-world deployment to achieve full-pose control accuracy comparable to the state-of-the-art centralized controller [6], and is deployed fully onboard. In addition, experiments with real MAVs demonstrate that our method remains robust under load model uncertainties, operates effectively in heterogeneous agent settings where one MAV uses a different controller, and remains functional even when one of the MAVs completely fails.
@@ 2 Related works
Cooperative aerial manipulation of a cable-suspended load typically embraces a centralized paradigm to consider the cable-load-MAVs system as a whole and requires global state observations to ensure safety and performance. Early research on multi-MAV cable-suspended load problems often relied on model simplifications, such as assuming a quasi-static regime to ignore dynamic coupling effects [12, 13, 14, 15], which cannot address force-related constraints and perform dynamic motions. Another class of methods leverages system flatness [16] and dynamic equations to account for dynamic coupling effects. An example is the cascaded scheme, which employs an outer-loop geometric controller to generate the commanded wrench for the load, distributes it as desired cable tensions, and executes it through inner-loop controllers of MAVs [3, 17, 18, 19]. The outer-loop controller can be replaced by various approaches, such as inverse dynamics control [20], linear quadratic regulator [4], and nonlinear model predictive control (NMPC) [5]. Recent work [6] leverages whole-body dynamics and NMPC to generate reference trajectories followed by an adaptive low-level controller, showing high agility and accuracy.
@@
However, these centralized methods require exponentially higher computational budgets and communication burdens with the number of agents involved. Therefore, decentralized controllers, such as distributed MPC [21, 22] have been proposed and tested in simulation to address the problem with the computational issues. But these methods still require reliable inter-agent data transfer to obtain real-time states from other agents, which does not fundamentally solve the problems with limited communication bandwidth.
@@
Multi-agent reinforcement learning has been extensively studied for complex multi-agent systems, including cooperative scenarios [23, 24, 25]. Beyond achieving expert-level performance in video games [26, 27], MARL has been successfully applied to robotics, enabling decentralized control of multiple agents. For instance, researchers have leveraged MARL to develop cooperative strategies in robot football [28, 29], as well as multi-robot object manipulation with quadrupedal robots, including pushing [30] and cable-based towing [31]. Unlike our approach, these manipulation methods [30, 31] rely on neighboring agent information through communication or onboard perception. In many cases, MARL is employed to optimize high-level task objectives while relying on mid- and low-level controllers for motor and sub-task execution, capitalizing on RL’s ability to optimize a long-horizon task-level objective [32].
@@
Recent work by [33] demonstrates MARL’s potential for cooperative object manipulation using simulated humanoids, relying solely on object bounding box information without explicit inter-agent communication. However, their approach depends on handcrafted reward functions that guide the humanoids toward predefined grasping points and walking behaviors. In MAV applications, MARL has been explored for tasks like swarming [34], but challenges remain due to the platform’s agility, instability, and reliance on high-frequency, low-latency control [35]. Recently, MARL has shown potential for training multi-MAV lifting systems using global state observations [36]. However, a significant challenge remains to address the sim-to-real gap and partial observability, especially for the multi-MAV lifting system, where dynamic uncertainties are substantial due to complex aerodynamic disturbances and unknown cable tensions.
@@ 3 Methods
As the problem is partially observable, we use a history of observations by stacking the current and last 2 observations of the policy [39]. For a more detailed discussion on the history length, we refer the readers to Appendix A.6.
We train the policies using MAPPO [8], a model-free MARL algorithm that extends PPO [40] with CTDE. The reward at time $t$, denoted as $r_{t}$, is defined as:
@@
Action space and low-level controller To balance reliable sim-to-real transfer with sufficient control authority, the choice of action space is critical. Prior work in single MAV control demonstrates that high-level outputs (e.g., position or velocity) enhance robustness to disturbances and sim-to-real gaps but limit performance, whereas low-level outputs (e.g., snap) improve tracking precision at the cost of larger transfer discrepancies [41, 42]. To address this trade-off, we propose a mid-level action space in desired accelerations and body rates (ACCBR). This approach preserves adequate control capability while also being robust against uncertain disturbances and model mismatches from the cable-suspended load.
@@
where external forces $\bm{f}_{\mathrm{ext}}$, primarily due to the cable tensions, are estimated using the MAV mass $\mathit{m}_{i}$, filtered accelerometer measurements $\bm{a}_{i,\mathrm{filtered}}$ and collective thrust $\bm{f}_{i,\mathrm{filtered}}$ computed from a classical quadratic thrust model and filtered rotor speed feedbacks [11]. The desired attitude command and the policy output body-rate command are then sent to the INDI attitude controller to generate rotor speed commands. We refer readers to [9, 10, 11] for further details on INDI.
Training setup We train our method completely in simulation and achieve zero-shot transfer to real-world experiments. The simulation environment is built using NVIDIA’s Isaac Lab [43], and the MARL algorithms are modified from [44]. Training was conducted on a consumer-grade RTX 3090 GPU and completed in 17 hours. The network architecture is a 4-layer MLP of size $[1024,512,256,128]$ for both the shared policies and the centralized critic. The inputs to the network are normalized stacked observation histories with history size $H=3$. For a complete overview of training details, network and agent parameters, we refer the readers to Appendix A.8.
@@ 4.1 Real-world experiments
Figure 3: Time series of pose tracking results comparing our method and a centralized NMPC method [6]. Our method also includes a setup with 4 MAVs.
Setpoint tracking Our real-world experiments demonstrate agile pose control of three MAVs with a cable-suspended load, tracking a 2 m displacement with (30°, -20°, -90°) attitude commands. We compare our decentralized method with the state-of-the-art centralized NMPC approach [6] in Figure 3. Despite being fully decentralized, our method achieves comparable tracking performance with positional and attitude RMSEs of 0.52 m (vs 0.45 m) and 22.93° (vs 16.24°), respectively. Note that RMSE comparisons favor NMPC as it tracks a reference trajectory while we only track target poses, resulting in a larger RMSE in the transient area of the step command. The time-to-target (error $\text{0.10}\,\text{m}/\text{10}^{\circ}$) is 6.84 s for NMPC vs. 8.36 s for ours, and the final displacement (RMSE) is 0.05 m /4.02 for NMPC vs. 0.04 m /5.78 for ours. We also show successful pose control with 4 MAVs (without cable slack), resulting in tracking RMSEs of 0.92 m and 42.67°. The increased error, compared to the 3 MAV case, may be due to the system becoming overconstrained, which introduces more complex coordination and (cable) dynamics [18]. In terms of computational efficiency, we run the NMPC and our method onboard a Raspberry Pi 5 (2.4 GHz quad-core ARM Cortex-A76). Our method inferences in 6 ms at 100 Hz, versus NMPC’s 78 ms at 10 Hz. Crucially, while NMPC’s computation time grows exponentially with agent count, e.g., 174 ms and 267 ms for 5 and 6 agents respectively, our agent-independent approach maintains constant computation time regardless of team size.
@@
Heterogeneous agents Although our policy is trained under the assumption of homogeneous agents, it remains effective when deployed with heterogeneous agents. In this experiment, we let the load hover at a fixed point. Then we hacked one of the MAVs by replacing its RL policy with a model-based controller [11], and provided it with different setpoints to observe the behavior of the other two MAVs controlled by the RL policy. Specifically, we commanded the hacked MAV to move outwards on the y-axis by 0.7 m to pull the load away from the reference; we then commanded the hacked MAV to move inwards by 0.3 m to push it closer to the other two MAVs. Figure 4A provides a snapshot of the experiments. Since the policy is conditioned solely on the load pose and not on the states of the other agents, the two remaining MAVs utilizing the policy can compensate for load pose deviations from the reference. In contrast, the fully observable policy fails in these conditions due to the dependence on the states of all agents. Time series can be found in Appendix A.2.
@@ 4.2 Comparison among different action and observation spaces
We compare our selected observation and action spaces with alternatives in simulation for safety. The Agilicious flight stack is used with the Gazebo simulator [45] and RotorS [46] plugins, which add sensor noise, aerodynamic disturbances, and system latencies in a ROS environment. All policies are trained for 1 billion environment steps (10 h) and evaluated 10 times in Gazebo.
@@
The RMSE results in Table 6 demonstrate that the VEL action space achieves the best performance, followed by ACCBR, while ACC fails to track the load orientation accurately. Notably, the widely used CTBR approach [47, 32] fails to learn effectively. Since CTBR directly commands collective thrust without leveraging the proposed low-level controller’s disturbance compensation, we hypothesize that the unpredictable cable forces exerted on each MAV make the learning process prohibitively difficult, as there are no cable force sensors mounted for both training and evaluations.
@@
Observation space To benchmark the decentralized policy’s performance, we compare three observation space cases: (1) the fully observable case with global state $\bm{s}=[\bm{x}_{L},\bm{x}_{G},\bm{x}_{1},\bm{x}_{2},\bm{x}_{3}]$, (2) an augmented partial observability case where each MAV $i$ also receives the load twist and other MAVs’ positions ("Partial augmented") $\bm{o}_{i}=[\bm{x}_{L},\bm{x}_{G},\bm{p}_{j_{1}},\bm{p}_{j_{2}},\bm{x}_{i},\bm{e}_{i}]$ with $\bm{p}_{j_{1}},\bm{p}_{j_{2}}$ representing the neighboring agents’ positions, and (3) the partially observable case. For partially observable cases, we include observation histories ($H=3$) to improve state estimation and decision-making under uncertainty [39]. Figure 7 reveals comparable convergence across all configurations, indicating that load pose alone serves as a sufficient statistic for implicit MAV coordination, while the full global state contains redundant elements.

### v3

@@ 1 Introduction
Autonomous Micro Aerial Vehicles (MAVs) offer great capability for transporting slung loads to dangerous and remote locations [barmpounakis_unmanned_2016]. While a single low-cost MAV has limited payload capacity, collaborative teams of MAVs can transport significantly heavier loads. In addition, by connecting each MAV with the load at different points using tethers, the full pose of the load can be controlled by changing the position of the MAVs, yielding a cooperative cable-suspended manipulation solution, which shows great potential for aerial-based construction, inspection, and resecuring [sreenath_dynamics_2013, lee_geometric_2018, geng_load-distribution-based_2022, li_nonlinear_2023, sun2025agilecooperativeaerialmanipulation].
@@
To coordinate and control MAV fleets, the state-of-the-art method [sun2025agilecooperativeaerialmanipulation] employs a centralized framework that accurately captures the strong dynamical coupling between the MAVs and the suspended load. This ensures safety and stability while addressing the significant underactuation inherent to cable-suspended systems, preventing actuator saturations and reciprocal collisions. However, using centralized control strategies for such systems suffers from critical drawbacks: computational complexity tends to scale exponentially with the number of agents for many approaches, rendering real-time control infeasible for larger teams with a centralized scheme [sun2025agilecooperativeaerialmanipulation, bakule_decentralized_2012]. In addition, dependence on global state information and centralized communication is often impractical due to limits on sensors and communication bandwidth. A plausible solution, decentralization, remains an open challenge to effectively coordinate MAV fleets due to partial observability, limited communication bandwidth, and decision-making under strong dynamical coupling between agents while co-manipulating an object.
@@
In this work, we present the first decentralized algorithm to achieve a real-world demonstrated full-pose manipulation of a cable-suspended payload using a team of MAVs. Our method leverages multi-agent reinforcement learning (MARL) and does not require any inter-agent communication. Instead, each agent only takes their own state and identity, the load pose, and the target load pose as observations. We train the policy through MARL in a centralized training with decentralized execution (CTDE) paradigm using multi-agent proximal policy optimization (MAPPO) [yu_surprising_2022]. Each MAV learns to communicate implicitly through the load pose information. To fill the sim-to-real gap in this highly dynamic cooperative task, we design the action space of the reinforcement learning (RL) policy as reference linear accelerations and body rates of the MAV and combine the RL policy with a low-level controller based on incremental nonlinear dynamic inversion (INDI) [smeur_adaptive_2016, tal_accurate_2021, sun_comparative_2022]. The low-level controller follows the linear acceleration command with the body rate reference as the feedforward commands, ensuring agile and smooth control maneuvers during the cooperative manipulation.
@@
Our method enables zero-shot transfer of the policy from simulation to real-world deployment to achieve full-pose control accuracy comparable to the state-of-the-art centralized controller [sun2025agilecooperativeaerialmanipulation], and is deployed fully onboard. In addition, experiments with real MAVs demonstrate that our method remains robust under load model uncertainties, operates effectively in heterogeneous agent settings where one MAV uses a different controller, and remains functional even when one of the MAVs completely fails.
@@ 2 Related works
Cooperative aerial manipulation of a cable-suspended load typically embraces a centralized paradigm to consider the cable-load-MAVs system as a whole and requires global state observations to ensure safety and performance. Early research on multi-MAV cable-suspended load problems often relied on model simplifications, such as assuming a quasi-static regime to ignore dynamic coupling effects [fink_planning_2011, michael_cooperative_2011, manubens_motion_2013, sanalitro_full-pose_2020], which cannot address force-related constraints and perform dynamic motions. Another class of methods leverages system flatness [sreenath_geometric_2013] and dynamic equations to account for dynamic coupling effects. An example is the cascaded scheme, which employs an outer-loop geometric controller to generate the commanded wrench for the load, distributes it as desired cable tensions, and executes it through inner-loop controllers of MAVs [lee_geometric_2018, li_cooperative_2021, li_rotortm_2024, wahba_efficient_2024]. The outer-loop controller can be replaced by various approaches, such as inverse dynamics control [masone_shared_2021], linear quadratic regulator [geng_load-distribution-based_2022], and nonlinear model predictive control (NMPC) [li_nonlinear_2023]. Recent work [sun2025agilecooperativeaerialmanipulation] leverages whole-body dynamics and NMPC to generate reference trajectories followed by an adaptive low-level controller, showing high agility and accuracy.
@@
However, these centralized methods require exponentially higher computational budgets and communication burdens with the number of agents involved. Therefore, decentralized controllers, such as distributed MPC [wehbeh_distributed_2020, wang_auto-multilift_2024] have been proposed and tested in simulation to address the problem with the computational issues. But these methods still require reliable inter-agent data transfer to obtain real-time states from other agents, which does not fundamentally solve the problems with limited communication bandwidth.
@@
Multi-agent reinforcement learning has been extensively studied for complex multi-agent systems, including cooperative scenarios [busoniu_comprehensive_2008, zhang_multi-agent_2021, sukthankar_cooperative_2017]. Beyond achieving expert-level performance in video games [vinyals_grandmaster_2019, openai_dota_2019], MARL has been successfully applied to robotics, enabling decentralized control of multiple agents. For instance, researchers have leveraged MARL to develop cooperative strategies in robot football [liu_motor_2022, li_marladona_2024], as well as multi-robot object manipulation with quadrupedal robots, including pushing [feng_learning_2024] and cable-based towing [chen2025decentralized]. Unlike our approach, these manipulation methods [feng_learning_2024, chen2025decentralized] rely on neighboring agent information through communication or onboard perception. In many cases, MARL is employed to optimize high-level task objectives while relying on mid- and low-level controllers for motor and sub-task execution, capitalizing on RL’s ability to optimize a long-horizon task-level objective [song_reaching_2023].
@@
Recent work by [gao_coohoi_2024] demonstrates MARL’s potential for cooperative object manipulation using simulated humanoids, relying solely on object bounding box information without explicit inter-agent communication. However, their approach depends on handcrafted reward functions that guide the humanoids toward predefined grasping points and walking behaviors. In MAV applications, MARL has been explored for tasks like swarming [batra_decentralized_2022], but challenges remain due to the platform’s agility, instability, and reliance on high-frequency, low-latency control [xingbootstrapping]. Recently, MARL has shown potential for training multi-MAV lifting systems using global state observations [xu_omnidrones_2024]. However, a significant challenge remains to address the sim-to-real gap and partial observability, especially for the multi-MAV lifting system, where dynamic uncertainties are substantial due to complex aerodynamic disturbances and unknown cable tensions.
@@ 3 Methods
As the problem is partially observable, we use a history of observations by stacking the current and last 2 observations of the policy [hausknecht2015deep]. For a more detailed discussion on the history length, we refer the readers to Appendix A.6.
We train the policies using MAPPO [yu_surprising_2022], a model-free MARL algorithm that extends PPO [schulman_proximal_2017] with CTDE. The reward at time $t$, denoted as $r_{t}$, is defined as:
@@
Action space and low-level controller To balance reliable sim-to-real transfer with sufficient control authority, the choice of action space is critical. Prior work in single MAV control demonstrates that high-level outputs (e.g., position or velocity) enhance robustness to disturbances and sim-to-real gaps but limit performance, whereas low-level outputs (e.g., snap) improve tracking precision at the cost of larger transfer discrepancies [kaufmann_benchmark_2022, eschmann_learning_2024]. To address this trade-off, we propose a mid-level action space in desired accelerations and body rates (ACCBR). This approach preserves adequate control capability while also being robust against uncertain disturbances and model mismatches from the cable-suspended load.
@@
where external forces $\bm{f}_{\mathrm{ext}}$, primarily due to the cable tensions, are estimated using the MAV mass $\mathit{m}_{i}$, filtered accelerometer measurements $\bm{a}_{i,\mathrm{filtered}}$ and collective thrust $\bm{f}_{i,\mathrm{filtered}}$ computed from a classical quadratic thrust model and filtered rotor speed feedbacks [sun_comparative_2022]. The desired attitude command and the policy output body-rate command are then sent to the INDI attitude controller to generate rotor speed commands. We refer readers to [smeur_adaptive_2016, tal_accurate_2021, sun_comparative_2022] for further details on INDI.
Training setup We train our method completely in simulation and achieve zero-shot transfer to real-world experiments. The simulation environment is built using NVIDIA’s Isaac Lab [mittal_orbit_2023], and the MARL algorithms are modified from [serrano2023skrl]. Training was conducted on a consumer-grade RTX 3090 GPU and completed in 17 hours. The network architecture is a 4-layer MLP of size $[1024,512,256,128]$ for both the shared policies and the centralized critic. The inputs to the network are normalized stacked observation histories with history size $H=3$. For a complete overview of training details, network and agent parameters, we refer the readers to Appendix A.8.
@@ 4.1 Real-world experiments
Figure 3: Time series of pose tracking results comparing our method and a centralized NMPC method [sun2025agilecooperativeaerialmanipulation]. Our method also includes a setup with 4 MAVs.
Setpoint tracking Our real-world experiments demonstrate agile pose control of three MAVs with a cable-suspended load, tracking a 2 m displacement with (30°, -20°, -90°) attitude commands. We compare our decentralized method with the state-of-the-art centralized NMPC approach [sun2025agilecooperativeaerialmanipulation] in Figure 3. Despite being fully decentralized, our method achieves comparable tracking performance with positional and attitude RMSEs of 0.52 m (vs 0.45 m) and 22.93° (vs 16.24°), respectively. Note that RMSE comparisons favor NMPC as it tracks a reference trajectory while we only track target poses, resulting in a larger RMSE in the transient area of the step command. The time-to-target (error $\text{0.10}\,\text{m}/\text{10}^{\circ}$) is 6.84 s for NMPC vs. 8.36 s for ours, and the final displacement (RMSE) is 0.05 m /4.02 for NMPC vs. 0.04 m /5.78 for ours. We also show successful pose control with 4 MAVs (without cable slack), resulting in tracking RMSEs of 0.92 m and 42.67°. The increased error, compared to the 3 MAV case, may be due to the system becoming overconstrained, which introduces more complex coordination and (cable) dynamics [li_rotortm_2024]. In terms of computational efficiency, we run the NMPC and our method onboard a Raspberry Pi 5 (2.4 GHz quad-core ARM Cortex-A76). Our method inferences in 6 ms at 100 Hz, versus NMPC’s 78 ms at 10 Hz. Crucially, while NMPC’s computation time grows exponentially with agent count, e.g., 174 ms and 267 ms for 5 and 6 agents respectively, our agent-independent approach maintains constant computation time regardless of team size.
@@
Heterogeneous agents Although our policy is trained under the assumption of homogeneous agents, it remains effective when deployed with heterogeneous agents. In this experiment, we let the load hover at a fixed point. Then we hacked one of the MAVs by replacing its RL policy with a model-based controller [sun_comparative_2022], and provided it with different setpoints to observe the behavior of the other two MAVs controlled by the RL policy. Specifically, we commanded the hacked MAV to move outwards on the y-axis by 0.7 m to pull the load away from the reference; we then commanded the hacked MAV to move inwards by 0.3 m to push it closer to the other two MAVs. Figure 4A provides a snapshot of the experiments. Since the policy is conditioned solely on the load pose and not on the states of the other agents, the two remaining MAVs utilizing the policy can compensate for load pose deviations from the reference. In contrast, the fully observable policy fails in these conditions due to the dependence on the states of all agents. Time series can be found in Appendix A.2.
@@ 4.2 Comparison among different action and observation spaces
We compare our selected observation and action spaces with alternatives in simulation for safety. The Agilicious flight stack is used with the Gazebo simulator [koenig2004design] and RotorS [Furrer2016] plugins, which add sensor noise, aerodynamic disturbances, and system latencies in a ROS environment. All policies are trained for 1 billion environment steps (10 h) and evaluated 10 times in Gazebo.
@@
The RMSE results in Table 6 demonstrate that the VEL action space achieves the best performance, followed by ACCBR, while ACC fails to track the load orientation accurately. Notably, the widely used CTBR approach [kaufmann_champion-level_2023, song_reaching_2023] fails to learn effectively. Since CTBR directly commands collective thrust without leveraging the proposed low-level controller’s disturbance compensation, we hypothesize that the unpredictable cable forces exerted on each MAV make the learning process prohibitively difficult, as there are no cable force sensors mounted for both training and evaluations.
@@
Observation space To benchmark the decentralized policy’s performance, we compare three observation space cases: (1) the fully observable case with global state $\bm{s}=[\bm{x}_{L},\bm{x}_{G},\bm{x}_{1},\bm{x}_{2},\bm{x}_{3}]$, (2) an augmented partial observability case where each MAV $i$ also receives the load twist and other MAVs’ positions ("Partial augmented") $\bm{o}_{i}=[\bm{x}_{L},\bm{x}_{G},\bm{p}_{j_{1}},\bm{p}_{j_{2}},\bm{x}_{i},\bm{e}_{i}]$ with $\bm{p}_{j_{1}},\bm{p}_{j_{2}}$ representing the neighboring agents’ positions, and (3) the partially observable case. For partially observable cases, we include observation histories ($H=3$) to improve state estimation and decision-making under uncertainty [hausknecht2015deep]. Figure 7 reveals comparable convergence across all configurations, indicating that load pose alone serves as a sufficient statistic for implicit MAV coordination, while the full global state contains redundant elements.